To test whether a greater proportion of Canadians than Americans believe Edward Snowden is a hero, a total of 1,519 Canadians and 1010 Americans were randomly selected. Out of the total number of Canadians chosen 1018 responded he is a hero against 515 Americans who claimed he is a hero. Calculate the difference in the sample proportion?

Select one: a. 0.67 b. 0.51 c. 0.16 d. 0.23
Answer #1

Solution:-
=> c. 0.16

Explanation :

the difference in the sample proportion : p1 - p2 = 0.67 - 0.51 = 0.16

p1 = x1/n1 = 1018/1519 = 0.67

p2 = x2/n2 = 515/1010 = 0.51